Find the angle between the vectors a and b given that

a = 3i + 4j and b = 5i – 12j (correct to 1d.p)


1
Expert's answer
2021-09-16T17:41:45-0400

"cos\theta=\frac{ab}{|a||b|}=\frac{3*5+4(-12)}{\sqrt{3^2+4^2}\sqrt{5^2+(-12)^2}}=\frac{-33}{5*13}=-0.5077."

"\theta=arccos(-0.5077)=120.5^0" .